# Alpha-Beta HMM: Hidden Markov Model Filtering with Equal Exit Probabilities and a Step-Size Parameter

### v2

@@ 1 Introduction
1.1 Proposed Filtering Algorithm
@@
where $\alpha\in\left(0,1/M\right)$ controls the weight assigned to the prior belief. Intuitively, $\alpha$ can be interpreted as the exit probability — the probability that the underlying state transitions to a different value between two consecutive time instants (see Sec. 1B for the formal definition). and $\beta>0$ is a step-size parameter. In the sequel we will demonstrate that this algorithm, referred to as $\alpha\beta$-HMM filtering, exhibits favorable learning performance in dynamic environments. We further elucidate the motivation behind this belief update rule, which can be interpreted as a simplified and generalized version of the hidden Markov model for online filtering.
1.2 Interpretation of $\alpha\beta$-HMM
@@ 1.4 Notation
1.5 Preview of Results
In this paper, we study the dynamical system underlying the evolution of the $\alpha\beta$-HMM and analyze its learning and adaptation performance. Due to the nonlinearity and stochasticity of the dynamical system corresponding to the $\alpha\beta$-HMM, we first introduce a deterministic reference dynamical system. Using the Contraction Mapping Theorem[31], we prove that in a steady state environment, the reference dynamical system convergences to a fixed point and derive the convergence rate (Theorem 2.4). This results allows us to analyze the adaptability of the $\alpha\beta$-HMM and its dependence on parameters including the exit probability $\alpha$, step-size parameter $\beta$ and the informativeness of observations. Subsequently, we calculate the position of the fixed point (Theorem 2.11) to analyze the learning performance of the $\alpha\beta$-HMM and its sensitivity to parameters. Finally, we prove that the distance between the original dynamical system and the reference system can be bounded, and by leveraging the fixed point of the reference system, we provide an upper bound for the instantaneous error probability of the algorithm (Theorem 3.16), thereby identifying factors that influence the error probability. The flow of the argument is summarized in Fig. 1.
@@ 1.5 Preview of Results
2 Steady State Dynamical Analysis
@@
2.1 Convergence Rate and Adaption Ability
@@
Proof 2.1.
Appendix 7.
@@
Proof 2.3.
Supplementary Material I.
Lemma 2.2 shows that the reference dynamical system has a fixed point, which is strictly negative for all components. To verify this, for all $m=1,\ldots,M-1$, when $01-\alpha M>$ 0 we have:
@@
Proof 2.5.
Appendix 8
Remark 2.6.
Theorem 2.4 provides an upper bound on the convergence rate of $\hat{x}_{i}$ to $\hat{x}^{\infty}$ in the sense of the $L_{\infty}$-norm, which decreases as the exit probability $\alpha$ increases, as shown in (43). We can also observe the impact of the data, captured in $\beta\underline{d}$, and the initial setting, captured in $\overline{x}_{0}$, on the convergence rate of the reference dynamical system to its fixed point. The parameter $\underline{d}$ depends on both the local likelihood function (e.g., pre-trained models) and the quality of observations (e.g., variance of local data noise). This parameter, referred to as the identifiability parameter, reflects the ability to distinguish between the two most challenging states. From Theorem 2.4, it follows that higher identifiability $\underline{d}$ and higher $\beta$ accelerate the convergence to the fixed point, while the deviation in the initial setting, captured in $\overline{x}_{0}$, slows down the convergence rate.
@@
Proof 2.8.
Supplementary Material II.
Corollary 2.7 provides a more precise upper bound on the convergence rate of the reference dynamical system to its fixed point. It shows that as $\alpha\to 1/M$, the contraction coefficient $\gamma\to 0$, indicating that the rate of convergence accelerates. This observation aligns with intuition: when $\alpha\to 1/M$ in equation (5), the system relies less on historical observations and becomes more reliant on the most recently observed data. In subsequent numerical experiments, we observe that, generally larger values of $\alpha<1/M$ lead to faster convergence. Additionally, as $\beta\to\infty$, $\gamma\to 0$, which shows that increasing the value of $\beta$ leads to a faster convergence rate of the system to its fixed point.
@@
Remark 2.9.
The convergence rate of the reference dynamical system to its fixed point, as presented in Theorem 2.4, plays a crucial role in understanding the adaptation capabilities of the algorithm. Specifically, when the environment remains in a fixed state for an extended period, the reference dynamical system converges to its fixed point at an exponential rate. When the environment state changes, the corresponding fixed point shifts accordingly, and the system will converge to the new fixed point exponentially fast, as illustrated in Fig. 2. The convergence speed depends on the choice of the parameters $\alpha$ and $\beta$, and the quality of the data and model. When using the traditional Bayesian update (6), as indicated by equations (17) and (27) with $\alpha=0$ and $\beta=1$, the corresponding deterministic dynamical system simplifies to the linear equation
@@
Following Theorem 2.4, we can see that:
@@
where $\hat{x}_{\text{s}1}^{\infty}<0$ and $\hat{x}_{\text{s}2}^{\infty}>0$ are the two fixed points before and after $T_{1}$, respectively, and $\lambda_{\text{s}1}$ and $\lambda_{\text{s}2}$ are the convergence coefficients in two stages defined as in Theorem 2.4.
@@
2.2 Fixed Point Value and Learning Performance
In this section, we quantify the learning performance of the $\alpha\beta$-HMM and examine its dependence on the exit probability $\alpha$, the step-size parameter $\beta$, and the identifiability parameter $\underline{d}$. Since we have already proven that the reference dynamical system converges to the unique fixed point as described in Lemma 2.2, the position of the fixed point will be sufficient to characterize the resulting learning performance. Due to the nonlinearity of (37), explicitly solving for the exact fixed point value is challenging. Nevertheless, we can provide an estimate of its value. In Theorem 2.11, we derive upper and lower bounds for the counterpart of the belief $\boldsymbol{\mu}_{i}(\cdot)$ in the reference dynamical system.
@@ 2.2 Fixed Point Value and Learning Performance
Proof 2.12.
Appendix 9
Remark 2.13.
The upper and lower bounds in Theorem 2.11, as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$, are illustrated in Figure 3. We observe that when $\alpha$ is smaller and $\beta\underline{d}$ is larger, the steady-state belief on the correct state in the reference dynamical system is higher, indicating improved learning performance. This observation aligns with intuition: better learning performance is expected in less volatile environments, with lower observation noise, and a more accurate local likelihood model. The step-size parameter $\beta$ and the identifiability parameter $\underline{d}$ play a decisive role; when either $\beta$ or $\underline{d}$ is large, good learning performance is ensured in deterministic dynamical system, noise-free by construction, even when $\alpha$ is relatively large. This means that, in the deterministic setting, accurate inference is possible in volatile environments, provided that the quality of the data model and observations is high enough or that the value of $\beta$ is large enough. However, choosing a large $\beta$ comes at the cost of amplifying the effect of observation noise, as we will further demonstrate in Section 3.
Figure 3: The visualizations of the upper and lower bounds given in Theorem 2.11 as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$. Top: Visualization of the lower bound function. The left and right subplots show how the lower bound varies with one parameter while the other is fixed, with $\alpha$ and $\beta\underline{d}$ set to specific values in each case. Bottom: Visualization of the upper bound function. Since the upper bound is a function of $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$ (where $\overline{d}>\underline{d}$), the three subplots illustrate how the upper bound varies with $\alpha$ and $\beta\overline{d}$ for different fixed values of $\beta\underline{d}$.
3 From Deterministic to Stochastic Systems
In this section, we discuss the relationship between the stochastic dynamical system induced by $\alpha\beta$-HMM, represented by (19), and the reference dynamical system (27), establishing that the dynamics of (27) are representative of the stochastic system (19), which describes the actual learning dynamics of the $\alpha\beta$-HMM recursion. In Lemma 3.14, we show that the expected distance between $\boldsymbol{x}_{i}$ and the fixed point of the reference dynamical system can be asymptotically bounded. Then, in Theorem 3.16, we compute the instantaneous error probability of the $\alpha\beta$-HMM based on the fixed point of the reference dynamical system. To ensure tractability of the theoretical analysis, we introduce the following assumption.
@@ 3 From Deterministic to Stochastic Systems
    where it is established in Theorem 2.4 that $\hat{x}_{m,i}\to\hat{x}^{\infty}_{m}$, as $i\to\infty$.
@@
Proof 3.15.
Appendix 10
In Lemma 3.14, we show that the distance between the original dynamical system and the fixed point of the reference dynamical system can be bounded. The coefficient $\lambda_{1}$ is always less than 1, then as the time instant approaches infinity, the upper bound of the expected distance between the two is related to the first absolute moment of the log-likelihood ratio of the observational model and the contraction coefficient $\lambda_{1}$. In other words, when the observation noise decreases, the stochastic dynamical system corresponding to the $\alpha\beta$-HMM (80) approaches the fixed point of the reference dynamical system.
@@
other variable definitions remain unchanged from those presented in Lemma 3.14. Furthermore, we have the following asymptotic estimate of the instantaneous error probability:
@@
Proof 3.17.
Appendix 11
Remark 3.18.
The contraction coefficient $\lambda_{1}$ in Theorem 3.16 is always less than 1, consequently the upper bound of the instantaneous error probability decays exponentially with time and eventually converges to a fixed value
@@
The asymptotic analysis results for small $\alpha$ further indicate that as $\alpha$ approaches zero, the instantaneous error probability diminishes over time. This phenomenon highlights the trade-off between the learning capability and the adaptation ability of the algorithm: while a small $\alpha$ yields good learning performance, it reduces the adaptation rate, as illustrated in (83). Moreover, when $\alpha=0$, the proposed algorithm degenerates into Bayes’ formula, which, as discussed in Remark 2.9, exhibits poor adaptability.
4 Parameter Selection and Tuning
Based on the analysis in Section 2, we summarize the impacts of the exit probability parameter $\alpha$ and the step-size parameter $\beta$ as follows:
@@ 4 Parameter Selection and Tuning
    The exit probability parameter $\alpha$ can be employed to adjust the algorithm’s adaptability and learning performance. In more volatile environments, a larger value of $\alpha$ is preferred to enhance adaptability (Theorem 2.4). Conversely, if the goal is to ensure superior learning performance, a smaller value of $\alpha$ should be selected (Theorems 2.11 and 3.16). This conclusion is consistent with intuition, since $\alpha$ reflects the volatility of the environment.
@@
    The step-size parameter $\beta$ serves to balance the algorithm’s adaptability, learning performance, and the influence of observation noise. Specifically, under the assumption of the identifiability of the underlying true state (Assumption 2), a larger $\beta$ results in a faster adaptation rate (Theorem 2.4, Corollary 2.7) and improved learning performance in noise-free scenarios (Theorem 2.11). However, since $\beta$ also amplifies the effect of noise (Theorem 3.16), selecting a large $\beta$ is beneficial when the observation noise is minimal, while in high-noise conditions a large $\beta$ may lead to instability in the algorithm’s performance. Notably, when $\beta$ is set to 1, the proposed $\alpha\beta$-HMM algorithm can be interpreted as the HMM model corresponding to the equal exit probability transition matrix.
5 Numerical Experiments
5.1 Algorithm performance and parameter tuning
@@
The influence of the step-size parameter $\beta$ under different levels of observation noise is illustrated in Fig. 5. In the case of less noisy observations ($\sigma^{2}=1$), a higher value of $\beta$ leads to improved final learning accuracy across all three parameter settings. In contrast, when the observation noise is larger ($\sigma^{2}=2$), the learning accuracy decreases as $\beta$ increases. These experimental results are consistent with the theoretical findings presented in Theorem 3.16.
@@
5.2 Trade-off between learning and adaption
The above results can be explained using the steady-state analysis results from Theorem 2.4 and Theorem 2.11. Assuming that $\theta_{i}^{\star}$ remains unchanged as $\theta_{0}$, i.e., $\theta_{i}^{\star}\equiv\theta^{\star}$, it is straightforward to compute:
@@ 5.2 Trade-off between learning and adaption
Thus, when $\theta^{\star}$ and $\sigma^{2}$ are fixed, the value of $d_{m}$ remains constant. In Fig. 6(a), we observe that increasing the value of $\alpha$ while keeping $\beta$ constant causes the fixed points of the reference dynamical system to approach zero, indicating lower confidence and higher sensitivity. At the same time, the convergence rate to the fixed point increases as $\alpha$ approaches $1/M$. Conversely, when $\alpha$ is fixed and $\beta$ is adjusted, larger values of $\beta$ lead to fixed points further from zero and faster convergence rates. These findings are consistent with the analytical results presented in Theorems 2.4 and 2.11.
@@
Figure 6: (a) Comparison of the log-belief ratio evolution in the reference dynamical system under different parameter settings. Before iteration $i=25$, the underlying true state is $\theta_{i}^{=}\theta_{0}$, and after that, $\theta_{i}^{=}\theta_{4}$. The fixed point, $\hat{x}^{\infty}$, is computed using (37) when $\theta_{i}^{\star}=\theta_{0}$. (b) Comparison of the true evolution of $\left\|\hat{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}$ and the evolution estimated using the contraction coefficient $\lambda$ in Theorem 2.4 on semilog y-axis coordinates. (c) Evolution of belief on the underlying true state in the reference dynamical system under different parameter settings. The estimated upper and lower bounds are computed using Theorem 2.11.
@@
In Fig. 6(b), we further compare the convergence rates of $\hat{x}_{m,i}$ to the fixed point under different parameter settings when $\theta_{i}^{\star}\equiv\theta_{0}$. The linear decay observed on the semilog Y-axis, the upper bound estimation, and the variation in convergence rate with respect to adjustments in parameters $\alpha$ and $\beta$ are all consistent with the conclusions of Theorem 2.4 and Corollary 2.7, which predict exponential convergence to the fixed point with a rate that increases with $\alpha$ and $\beta$. In Fig.6(c), we demonstrate the evolution of the belief on the underlying true state $\theta_{0}$ in the reference dynamical system, i.e., $\hat{\mu}_{i}(\theta_{0})$, under different parameter settings. It can be observed that as $\alpha$ increases and $\beta$ decreases, both the steady-state belief on the true state and its estimated lower bound gradually decline. Furthermore, as $\alpha$ and $\beta$ increase, the belief converges to the steady state more rapidly. These results align with the analysis in Theorem 2.4 and Theorem 2.11, emphasizing the trade-off between learning performance and adaptation capability through the tuning of $\alpha$, and highlighting the role of the step-size parameter $\beta$ in shaping the convergence behavior.
5.3 Error Probability of the algorithm

### v3

@@ I Introduction
I-A Proposed Filtering Algorithm
@@
where $\alpha\in\left(0,1/M\right)$ controls the weight assigned to the prior belief. Intuitively, $\alpha$ can be interpreted as the exit probability — the probability that the underlying state transitions to a different value between two consecutive time instants (see Sec. IB for the formal definition). and $\beta>0$ is a step-size parameter. In the sequel we will demonstrate that this algorithm, referred to as $\alpha\beta$-HMM filtering, exhibits favorable learning performance in dynamic environments. We further elucidate the motivation behind this belief update rule, which can be interpreted as a simplified and generalized version of the hidden Markov model for online filtering.
I-B Interpretation of $\alpha\beta$-HMM
@@ I-D Notation
I-E Preview of Results
In this paper, we study the dynamical system underlying the evolution of the $\alpha\beta$-HMM and analyze its learning and adaptation performance. Due to the nonlinearity and stochasticity of the dynamical system corresponding to the $\alpha\beta$-HMM, we first introduce a deterministic reference dynamical system. Using the Contraction Mapping Theorem[31], we prove that in a steady state environment, the reference dynamical system convergences to a fixed point and derive the convergence rate (Theorem 1). This results allows us to analyze the adaptability of the $\alpha\beta$-HMM and its dependence on parameters including the exit probability $\alpha$, step-size parameter $\beta$ and the informativeness of observations. Subsequently, we calculate the position of the fixed point (Theorem 2) to analyze the learning performance of the $\alpha\beta$-HMM and its sensitivity to parameters. Finally, we prove that the distance between the original dynamical system and the reference system can be bounded, and by leveraging the fixed point of the reference system, we provide an upper bound for the instantaneous error probability of the algorithm (Theorem 3), thereby identifying factors that influence the error probability. The flow of the argument is summarized in Fig. 1.
@@ I-E Preview of Results
II Steady State Dynamical Analysis
@@
II-A Convergence Rate and Adaption Ability
@@
Proof:
See Supplementary Material I due to page limitations. ∎
@@
Proof:
See Supplementary Material II due to page limitations. ∎
Lemma 2 shows that the reference dynamical system has a fixed point, which is strictly negative for all components. To verify this, for all $m=1,\ldots,M-1$, when $01-\alpha M>$ 0 we have:
@@
Proof:
See Supplementary Material III due to page limitations. ∎
Remark 2
Theorem 1 provides an upper bound on the convergence rate of $\hat{x}_{i}$ to $\hat{x}^{\infty}$ in the sense of the $L_{\infty}$-norm, which decreases as the exit probability $\alpha$ increases, as shown in (43). We can also observe the impact of the data, captured in $\beta\underline{d}$, and the initial setting, captured in $\overline{x}_{0}$, on the convergence rate of the reference dynamical system to its fixed point. The parameter $\underline{d}$ depends on both the local likelihood function (e.g., pre-trained models) and the quality of observations (e.g., variance of local data noise). This parameter, referred to as the identifiability parameter, reflects the ability to distinguish between the two most challenging states. From Theorem 1, it follows that higher identifiability $\underline{d}$ and higher $\beta$ accelerate the convergence to the fixed point, while the deviation in the initial setting, captured in $\overline{x}_{0}$, slows down the convergence rate.
@@
Proof:
See Supplementary Material IV due to page limitations. ∎
Corollary 1 provides a more precise upper bound on the convergence rate of the reference dynamical system to its fixed point. It shows that as $\alpha\to 1/M$, the contraction coefficient $\gamma\to 0$, indicating that the rate of convergence accelerates. This observation aligns with intuition: when $\alpha\to 1/M$ in equation (5), the system relies less on historical observations and becomes more reliant on the most recently observed data. In subsequent numerical experiments, we observe that, generally larger values of $\alpha<1/M$ lead to faster convergence. Additionally, as $\beta\to\infty$, $\gamma\to 0$, which shows that increasing the value of $\beta$ leads to a faster convergence rate of the system to its fixed point.
@@
Remark 3
The convergence rate of the reference dynamical system to its fixed point, as presented in Theorem 1, plays a crucial role in understanding the adaptation capabilities of the algorithm. Specifically, when the environment remains in a fixed state for an extended period, the reference dynamical system converges to its fixed point at an exponential rate. When the environment state changes, the corresponding fixed point shifts accordingly, and the system will converge to the new fixed point exponentially fast, as illustrated in Fig. 2. The convergence speed depends on the choice of the parameters $\alpha$ and $\beta$, and the quality of the data and model. When using the traditional Bayesian update (6), as indicated by equations (17) and (27) with $\alpha=0$ and $\beta=1$, the corresponding deterministic dynamical system simplifies to the linear equation
@@
Following Theorem 1, we can see that:
@@
where $\hat{x}_{\text{s}1}^{\infty}<0$ and $\hat{x}_{\text{s}2}^{\infty}>0$ are the two fixed points before and after $T_{1}$, respectively, and $\lambda_{\text{s}1}$ and $\lambda_{\text{s}2}$ are the convergence coefficients in two stages defined as in Theorem 1.
@@
II-B Fixed Point Value and Learning Performance
In this section, we quantify the learning performance of the $\alpha\beta$-HMM and examine its dependence on the exit probability $\alpha$, the step-size parameter $\beta$, and the identifiability parameter $\underline{d}$. Since we have already proven that the reference dynamical system converges to the unique fixed point as described in Lemma 2, the position of the fixed point will be sufficient to characterize the resulting learning performance. Due to the nonlinearity of (37), explicitly solving for the exact fixed point value is challenging. Nevertheless, we can provide an estimate of its value. In Theorem 2, we derive upper and lower bounds for the counterpart of the belief $\boldsymbol{\mu}_{i}(\cdot)$ in the reference dynamical system.
@@ II-B Fixed Point Value and Learning Performance
Proof:
See Supplementary Material V due to page limitations. ∎
Remark 4
The upper and lower bounds in Theorem 2, as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$, are illustrated in Figure 3. We observe that when $\alpha$ is smaller and $\beta\underline{d}$ is larger, the steady-state belief on the correct state in the reference dynamical system is higher, indicating improved learning performance. This observation aligns with intuition: better learning performance is expected in less volatile environments, with lower observation noise, and a more accurate local likelihood model. The step-size parameter $\beta$ and the identifiability parameter $\underline{d}$ play a decisive role; when either $\beta$ or $\underline{d}$ is large, good learning performance is ensured in deterministic dynamical system, noise-free by construction, even when $\alpha$ is relatively large. This means that, in the deterministic setting, accurate inference is possible in volatile environments, provided that the quality of the data model and observations is high enough or that the value of $\beta$ is large enough. However, choosing a large $\beta$ comes at the cost of amplifying the effect of observation noise, as we will further demonstrate in Section III.
Figure 3: The visualizations of the upper and lower bounds given in Theorem 2 as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$. Top: Visualization of the lower bound function. The left and right subplots show how the lower bound varies with one parameter while the other is fixed, with $\alpha$ and $\beta\underline{d}$ set to specific values in each case. Bottom: Visualization of the upper bound function. Since the upper bound is a function of $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$ (where $\overline{d}>\underline{d}$), the three subplots illustrate how the upper bound varies with $\alpha$ and $\beta\overline{d}$ for different fixed values of $\beta\underline{d}$.
III From Deterministic to Stochastic Systems
In this section, we discuss the relationship between the stochastic dynamical system induced by $\alpha\beta$-HMM, represented by (19), and the reference dynamical system (27), establishing that the dynamics of (27) are representative of the stochastic system (19), which describes the actual learning dynamics of the $\alpha\beta$-HMM recursion. In Lemma 3, we show that the expected distance between $\boldsymbol{x}_{i}$ and the fixed point of the reference dynamical system can be asymptotically bounded. Then, in Theorem 3, we compute the instantaneous error probability of the $\alpha\beta$-HMM based on the fixed point of the reference dynamical system. To ensure tractability of the theoretical analysis, we introduce the following assumption.
@@ III From Deterministic to Stochastic Systems
    where it is established in Theorem 1 that $\hat{x}_{m,i}\to\hat{x}^{\infty}_{m}$, as $i\to\infty$.
@@
Proof:
See Supplementary Material VI due to page limitations. ∎
In Lemma 3, we show that the distance between the original dynamical system and the fixed point of the reference dynamical system can be bounded. The coefficient $\lambda_{1}$ is always less than 1, then as the time instant approaches infinity, the upper bound of the expected distance between the two is related to the first absolute moment of the log-likelihood ratio of the observational model and the contraction coefficient $\lambda_{1}$. In other words, when the observation noise decreases, the stochastic dynamical system corresponding to the $\alpha\beta$-HMM (80) approaches the fixed point of the reference dynamical system.
@@
other variable definitions remain unchanged from those presented in Lemma 3. Furthermore, we have the following asymptotic estimate of the instantaneous error probability:
@@
Proof:
See Supplementary Material VII due to page limitations. ∎
Remark 5
The contraction coefficient $\lambda_{1}$ in Theorem 3 is always less than 1, consequently the upper bound of the instantaneous error probability decays exponentially with time and eventually converges to a fixed value
@@
The asymptotic analysis results for small $\alpha$ further indicate that as $\alpha$ approaches zero, the instantaneous error probability diminishes over time. This phenomenon highlights the trade-off between the learning capability and the adaptation ability of the algorithm: while a small $\alpha$ yields good learning performance, it reduces the adaptation rate, as illustrated in (83). Moreover, when $\alpha=0$, the proposed algorithm degenerates into Bayes’ formula, which, as discussed in Remark 3, exhibits poor adaptability.
IV Parameter Selection and Tuning
Based on the analysis in Section II, we summarize the impacts of the exit probability parameter $\alpha$ and the step-size parameter $\beta$ as follows:
@@ IV Parameter Selection and Tuning
    The exit probability parameter $\alpha$ can be employed to adjust the algorithm’s adaptability and learning performance. In more volatile environments, a larger value of $\alpha$ is preferred to enhance adaptability (Theorem 1). Conversely, if the goal is to ensure superior learning performance, a smaller value of $\alpha$ should be selected (Theorems 2 and 3). This conclusion is consistent with intuition, since $\alpha$ reflects the volatility of the environment.
@@
    The step-size parameter $\beta$ serves to balance the algorithm’s adaptability, learning performance, and the influence of observation noise. Specifically, under the assumption of the identifiability of the underlying true state (Assumption 2), a larger $\beta$ results in a faster adaptation rate (Theorem 1, Corollary 1) and improved learning performance in noise-free scenarios (Theorem 2). However, since $\beta$ also amplifies the effect of noise (Theorem 3), selecting a large $\beta$ is beneficial when the observation noise is minimal, while in high-noise conditions a large $\beta$ may lead to instability in the algorithm’s performance. Notably, when $\beta$ is set to 1, the proposed $\alpha\beta$-HMM algorithm can be interpreted as the HMM model corresponding to the equal exit probability transition matrix.
V Numerical Experiments
V-A Algorithm performance and parameter tuning
@@
The influence of the step-size parameter $\beta$ under different levels of observation noise is illustrated in Fig. 5. In the case of less noisy observations ($\sigma^{2}=1$), a higher value of $\beta$ leads to improved final learning accuracy across all three parameter settings. In contrast, when the observation noise is larger ($\sigma^{2}=2$), the learning accuracy decreases as $\beta$ increases. These experimental results are consistent with the theoretical findings presented in Theorem 3.
@@
V-B Trade-off between learning and adaption
The above results can be explained using the steady-state analysis results from Theorem 1 and Theorem 2. Assuming that $\theta_{i}^{\star}$ remains unchanged as $\theta_{0}$, i.e., $\theta_{i}^{\star}\equiv\theta^{\star}$, it is straightforward to compute:
@@ V-B Trade-off between learning and adaption
Thus, when $\theta^{\star}$ and $\sigma^{2}$ are fixed, the value of $d_{m}$ remains constant. In Fig. 6(a), we observe that increasing the value of $\alpha$ while keeping $\beta$ constant causes the fixed points of the reference dynamical system to approach zero, indicating lower confidence and higher sensitivity. At the same time, the convergence rate to the fixed point increases as $\alpha$ approaches $1/M$. Conversely, when $\alpha$ is fixed and $\beta$ is adjusted, larger values of $\beta$ lead to fixed points further from zero and faster convergence rates. These findings are consistent with the analytical results presented in Theorems 1 and 2.
@@
Figure 6: (a) Comparison of the log-belief ratio evolution in the reference dynamical system under different parameter settings. Before iteration $i=25$, the underlying true state is $\theta_{i}^{=}\theta_{0}$, and after that, $\theta_{i}^{=}\theta_{4}$. The fixed point, $\hat{x}^{\infty}$, is computed using (37) when $\theta_{i}^{\star}=\theta_{0}$. (b) Comparison of the true evolution of $\left\|\hat{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}$ and the evolution estimated using the contraction coefficient $\lambda$ in Theorem 1 on semilog y-axis coordinates. (c) Evolution of belief on the underlying true state in the reference dynamical system under different parameter settings. The estimated upper and lower bounds are computed using Theorem 2.
@@
In Fig. 6(b), we further compare the convergence rates of $\hat{x}_{m,i}$ to the fixed point under different parameter settings when $\theta_{i}^{\star}\equiv\theta_{0}$. The linear decay observed on the semilog Y-axis, the upper bound estimation, and the variation in convergence rate with respect to adjustments in parameters $\alpha$ and $\beta$ are all consistent with the conclusions of Theorem 1 and Corollary 1, which predict exponential convergence to the fixed point with a rate that increases with $\alpha$ and $\beta$. In Fig.6(c), we demonstrate the evolution of the belief on the underlying true state $\theta_{0}$ in the reference dynamical system, i.e., $\hat{\mu}_{i}(\theta_{0})$, under different parameter settings. It can be observed that as $\alpha$ increases and $\beta$ decreases, both the steady-state belief on the true state and its estimated lower bound gradually decline. Furthermore, as $\alpha$ and $\beta$ increase, the belief converges to the steady state more rapidly. These results align with the analysis in Theorem 1 and Theorem 2, emphasizing the trade-off between learning performance and adaptation capability through the tuning of $\alpha$, and highlighting the role of the step-size parameter $\beta$ in shaping the convergence behavior.
V-C Error Probability of the algorithm